# Do We Need Reformer for Vision? An Experimental Comparison with Vision Transformers

### v1

@@ 2 Related Works
In vision, self-attention can be applied by treating pixels or patches as tokens, enabling each location to aggregate information from other spatial locations through learned similarity scores. Because images have inherent spatial structure, positional encoding is essential to avoid permutation ambiguity. [cordonnier2020relationshipselfattentionconvolutionallayers] show that, with relative positional encoding, multi-head self-attention can simulate convolutional behavior. In particular, their Theorem 1 states that a multi-head self-attention layer with $N_{h}$ heads and a relative positional encoding of dimension $D_{p}\geq 3$ can express any convolutional layer with kernel size $\sqrt{N_{h}}\times\sqrt{N_{h}}$ and up to $\min(D_{h},D_{out})$ output channels. This result helps justify why attention-based vision backbones can inherit strong local inductive biases while still allowing more global interactions when needed, supporting our exploration of LSH-based attention as an efficiency-oriented alternative. To mitigate the quadratic cost of global attention, several vision transformers incorporate locality or hierarchical execution, methods that were firstly introduced in NLP. Swin-style models and their extensions introduce structured attention and improved token handling, aiming to reduce computation while maintaining competitive accuracy [9874052, 10643534]. Alternative attention formulations also explore replacing full quadratic interactions with structured approximations. Approaches such as attention via convolutional nearest neighbors highlight that attention-like behavior can be achieved through more constrained or efficient operators, potentially improving scalability depending on the target regime [kang2025attentionconvolutionalnearestneighbors].Recent survey work on Vision Transformers highlights that deploying ViTs efficiently on edge devices requires aggressive reductions in attention complexity, motivating alternative attention mechanisms such as structured, approximate, or hardware-friendly formulations. [saha2025visiontransformersedge] A recent benchmark study compares over 30 efficient Vision Transformers under identical training and evaluation settings, revealing that ViT often remains Pareto-optimal across accuracy, speed, and memory trade-offs. Their results emphasize how architectural strategies—such as sparse attention, token reduction, and hybrid attention—affect practical efficiency, supporting our exploration of Reformer-style approximations.[nauen2024whichtransformertofavor]
@@ 3.3 Experimental Design and Fair Comparison
  • Small-scale setting: CIFAR-10, a 32 x 32 pixel color image dataset wit 10 classes where the token sequence is relatively short.
@@ 3.4 CIFAR-10
Both models are trained for 50 epochs using AdamW with weight decay for regularization to prevent underfitting. We apply Random and Horizontal Flip augmentation to the dataset, as well as Random Crop to ensure the models properly fits the data. We employ a learning rate schedule with 5 warmup epochs followed by cosine annealing to stabilize early optimization and ensure consistent training dynamics across the Reformer and ViT baselines.
@@ 4.1 CIFAR-10: Short-Sequence Regime
(a) Locality sensitive hashing
@@
Table 1: Metrics
@@
4.2 Image-Net-100: Mid-Scale Setting
On ImageNet-100, the gap between the two architectures becomes more apparent. The ViT baseline reaches a top-1 accuracy of 76.70%, while the Vision Reformer attains 74.20%, yielding a modest but consistent advantage for the ViT. The difference is more pronounced in terms of efficiency: the ViT requires on average 363 s per epoch, compared to 517 s for the Reformer (see Figure 2(b)). This behavior reflects the fact that ImageNet-100 still operates in a moderate token regime (256 tokens), where dense self-attention remains highly optimized by existing GPU kernels. In contrast, the Reformer incurs additional overhead from hashing, sorting, and bucketed attention, without yet benefiting from the asymptotic gains of $\mathcal{O}(n\log n)$ attention. As a result, the Reformer cannot close the runtime gap and fails to translate its theoretical efficiency advantage into practical speedups at this scale.
(a) Locality sensitive hashing
@@ 4.2 Image-Net-100: Mid-Scale Setting
Figure 2: average time per epoch on ImageNet-100
@@ 4.3 Diabetic Retinopathy: High-Resolution, Longer Tokens
On the other hand, the ViT clearly outperforms the Reformer in predictive performance. The ViT achieves a top-1 accuracy of 74.56%, whereas the Reformer reaches 70.96%. This gap is consistent with the class imbalance of the dataset and is further reflected in the macro-precision and macro-recall scores reported in Table LABEL:dr_metrics. The Reformer suffers more from the imbalance, yielding lower recall on the minority disease classes, while the ViT is better able to maintain balanced performance across the five severity levels. Overall, these results suggest that, in this high-resolution medical setting, the Reformer begins to approach the ViT in efficiency but still trails in accuracy, particularly in terms of sensitivity to underrepresented classes.
Table 2: Metrics
@@ 5 Discussion
Across all three datasets, the ViT baseline outperforms the Vision Reformer in terms of classification accuracy, and it is also more efficient in the short- and mid-sequence regimes. On CIFAR-10 and ImageNet-100, where each image is represented by 256 tokens, dense self-attention remains highly optimized on current GPU hardware, and the additional overhead introduced by LSH attention makes the Reformer slower in practice. These results suggest that, for typical 2D image classification tasks with moderate token counts, standard ViTs are still a strong default choice.
@@
Overall, our study suggests that Reformer-style attention is not yet a drop-in replacement for standard ViT attention in 2D image classification with typical resolutions. However, the partial efficiency gains observed in the high-resolution medical setting indicate that LSH-based architectures could become more attractive in scenarios where sequence lengths are much higher than in our experiments. A natural direction is to consider modalities where the number of tokens can approach those of long-context NLP tasks, such as 3D medical volumes, video classification, or multi-frame temporal models, where each spatial–temporal patch contributes to a much longer sequence. In such settings, the Reformer’s asymptotic advantages may be more pronounced, and a careful exploration of hybrid designs—combining convolutional stems, hierarchical tokenization, and LSH attention at selected stages—could yield more favourable accuracy–efficiency trade-offs.
@@ 6 Conclusion
In this work, we investigated the Reformer architecture as an efficient attention backbone for vision and compared it to a standard ViT under aligned settings on CIFAR-10, ImageNet-100, and a high-resolution diabetic retinopathy dataset. Our experiments show that, in all three cases, the ViT achieves higher accuracy, and it remains faster in the short- and mid-sequence regimes. Only in the high-resolution medical setting, with 784 tokens per image, does the Reformer begin to close the efficiency gap, but the gain is modest and does not compensate for the loss in accuracy.

### v2

@@ 2 Related Works
In vision, self-attention can be applied by treating pixels or patches as tokens, enabling each location to aggregate information from other spatial locations through learned similarity scores. Because images have inherent spatial structure, positional encoding is essential to avoid permutation ambiguity. [cordonnier2020relationshipselfattentionconvolutionallayers] show that, with relative positional encoding, multi-head self-attention can simulate convolutional behavior. In particular, their Theorem 1 states that a multi-head self-attention layer with $N_{h}$ heads and a relative positional encoding of dimension $D_{p}\geq 3$ can express any convolutional layer with kernel size $\sqrt{N_{h}}\times\sqrt{N_{h}}$ and up to $\min(D_{h},D_{out})$ output channels. This result helps justify why attention-based vision backbones can inherit strong local inductive biases while still allowing more global interactions when needed, supporting our exploration of LSH-based attention as an efficiency-oriented alternative.
To mitigate the quadratic cost of global attention, several vision transformers incorporate locality or hierarchical execution, methods that were firstly introduced in NLP. Swin-style models and their extensions introduce structured attention and improved token handling, aiming to reduce computation while maintaining competitive accuracy [9874052, 10643534]. Alternative attention formulations also explore replacing full quadratic interactions with structured approximations. Approaches such as attention via convolutional nearest neighbors highlight that attention-like behavior can be achieved through more constrained or efficient operators, potentially improving scalability depending on the target regime [kang2025attentionconvolutionalnearestneighbors]. Recent survey work on Vision Transformers highlights that deploying ViTs efficiently on edge devices requires aggressive reductions in attention complexity, motivating alternative attention mechanisms such as structured, approximate, or hardware-friendly formulations [saha2025visiontransformersedge]. A recent benchmark study compares over 30 efficient Vision Transformers under identical training and evaluation settings, revealing that ViT often remains Pareto-optimal across accuracy, speed, and memory trade-offs, while also emphasizing how architectural strategies such as sparse attention, token reduction, and hybrid attention affect practical efficiency [nauen2024whichtransformertofavor].
Beyond generic efficiency benchmarks, recent work has explored domain-specific and system-level optimizations. Mobile U-ViT introduces a lightweight U-shaped Vision Transformer tailored for medical image segmentation, combining large-kernel CNN-based hierarchical embeddings with a shallow transformer bottleneck to balance local inductive bias, long-range modeling, and strict computational constraints [tang2025mobileuvit]. In parallel, TAFP-ViT targets efficiency from a hardware–software co-design perspective, proposing QKV computational fusion, adaptive token pruning, and specialized accelerator architectures to drastically reduce attention overhead and improve throughput and energy efficiency [xu2025tafpvit]. Together, these works highlight complementary directions to attention approximation—namely domain-aware architectural design and system-level optimization—reinforcing the relevance of exploring Reformer-style LSH attention as a model-level efficiency mechanism that remains orthogonal to hardware or task-specific assumptions.
@@ 3.3 Experimental Design and Fair Comparison
  • Small-scale setting: CIFAR-10, a 32 x 32 pixel color image dataset with 10 classes where the token sequence is relatively short.
@@ 3.4 CIFAR-10
Both models are trained for 50 epochs using AdamW with weight decay for regularization to prevent underfitting. We apply Random and Horizontal Flip augmentation to the dataset, as well as Random Crop to ensure the models properly fit the data. We employ a learning rate schedule with 5 warmup epochs followed by cosine annealing to stabilize early optimization and ensure consistent training dynamics across the Reformer and ViT baselines.
@@ 4.1 CIFAR-10: Short-Sequence Regime
(a) Locality sensitive hashing attention
@@
Table 1: Vision Transformers and Reformer architectures metrics.
@@
4.2 ImageNet-100: Mid-Scale Setting
On ImageNet-100, the gap between the two architectures becomes more apparent. The ViT baseline reaches a top-1 accuracy of 76.70%, while the Vision Reformer attains 74.20%, yielding a modest but consistent advantage for the ViT. The difference is more pronounced in terms of efficiency: the ViT requires on average 363 s per epoch, compared to 517 s for the Reformer (see Figure 2). This behavior reflects the fact that ImageNet-100 still operates in a moderate token regime (256 tokens), where dense self-attention remains highly optimized by existing GPU kernels. In contrast, the Reformer incurs additional overhead from hashing, sorting, and bucketed attention, without yet benefiting from the asymptotic gains of $\mathcal{O}(n\log n)$ attention. As a result, the Reformer cannot close the runtime gap and fails to translate its theoretical efficiency advantage into practical speedups at this scale.
(a) Locality-sensitive hashing attention
@@ 4.2 ImageNet-100: Mid-Scale Setting
Figure 2: average time per epoch on ImageNet-100 in seconds
@@ 4.3 Diabetic Retinopathy: High-Resolution, Longer Tokens
On the other hand, the ViT clearly outperforms the Reformer in predictive performance. The ViT achieves a top-1 accuracy of 74.56%, whereas the Reformer reaches 70.96%. This gap is consistent with the class imbalance of the dataset and is further reflected in the macro-precision and macro-recall scores reported in Table LABEL:TABLE_2. The Reformer suffers more from the imbalance, yielding lower recall on the minority disease classes, while the ViT is better able to maintain balanced performance across the five severity levels. Overall, these results suggest that, in this high-resolution medical setting, the Reformer begins to approach the ViT in efficiency but still trails in accuracy, particularly in terms of sensitivity to underrepresented classes.
Table 2: Vision Transformers and Reformer architectures metrics.
@@ 5 Discussion
Across all three datasets, the ViT baseline outperforms the Vision Reformer in terms of efficiency in the short- and mid-sequence regimes. On CIFAR-10 and ImageNet-100, where each image is represented by 256 tokens, dense self-attention remains highly optimized on current GPU hardware, and the additional overhead introduced by LSH attention makes the Reformer slower in practice. These results suggest that, for typical 2D image classification tasks with moderate token counts, standard ViTs are still a strong default choice.
@@
Overall, our study suggests that Reformer-style attention is not yet a drop-in replacement for standard ViT attention in 2D image classification with typical resolutions. However, the partial efficiency gains observed in the high-resolution medical setting indicate that LSH-based architectures could become more attractive in scenarios where sequence lengths are much higher than in our experiments. A natural direction is to consider modalities where the number of tokens can approach those of long-context NLP tasks, such as 3D medical volumes, video classification, or multi-frame temporal models, where each spatial–temporal patch contributes to a much longer sequence. In such settings, the Reformer’s asymptotic advantages may be more pronounced, and careful exploration of hybrid designs—combining convolutional stems, hierarchical tokenization, and LSH attention at selected stages—could yield more favorable accuracy–efficiency trade-offs.
@@ 6 Conclusion
In this work, we investigated the Reformer architecture as an efficient attention backbone for vision and compared it to a standard ViT under aligned settings on CIFAR-10, ImageNet-100, and a high-resolution diabetic retinopathy dataset. Our experiments show that, in all three cases, the ViT achieves higher accuracy except for CIFAR-10, and it remains faster in the short and mid-sequence regimes. Only in the high-resolution medical setting, with 784 tokens per image, does the Reformer begin to close the efficiency gap, but the gain is modest and does not compensate for the loss in accuracy.